OpenStudy (anonymous):

Dr.drake is thinkinking about retirment and decides to sail around the world once he retires. He buys a sailboat for 125000. He borrows the money at an arp of 7.5% for five years. WHat will his totoal intrest be?

OpenStudy (ash2326):

Interest=P*R*T P=125000 R=7.5% T=5 Interest=125000*7.5*5/100 =46875